OpenStudy (anonymous):

Find the 9th term of the geometric sequence with a2 = 0.3 and a6 = 0.00003

OpenStudy (lgbasallote):

an = a1r^n-1 (((0.3r)r)r) = 0.00003 0.3r^3 = 0.00003 r^3 = 0.0001 r = 0.0464158883 so a1 = 6.46330408 so a9 = (6.46330408)(0.0464158883)^8 a9 = 1.39 × 10^-10